### v1

# Examining Input Modalities and Visual Feedback Designs in Mobile Expressive Writing

Abstract.
Expressive writing is an established approach for stress management, and recent practices include information technology. Although mobile interfaces have the potential to support daily stress management practices, interface designs for such mobile expressive writing and their effects on stress relief still lack empirical understanding. To fill the gap, we examined the interface design of mobile expressive writing by investigating the influence of input modalities and visual feedback designs on usability and perceived cathartic effects through in-the-wild studies. While our studies confirmed the stress relief effects of mobile expressive writing, our results offer important insights in interface design. We found keyboard-based text entry more user-friendly and preferred over voice messages due to its privacy friendliness and reflection process. Participants expressed different reasons for preferring different post-writing visual feedback depending on the cause and type of stress. This paper also discusses future research opportunities in interface designs for mobile expressive writing.
1. Introduction
Excessive mental stress has a negative impact on many aspects of life, including mental and physical well-being (Veit and Ware, 1983; Hammen, 2005; Mushtaq et al., 2018). Due to its high global prevalence, several approaches to lower stress have been developed, primarily in the field of psychology. One established method is expressive writing, referring to the disclosure of distressing experiences and their related emotions typically by journaling (Pennebaker, 1997; Pennebaker and Chung, 2007). Expressive writing can provide people with a sense of stress relief (Pennebaker and Chung, 2007), known as a cathartic effect (Breuer and Freud, 1955). Expressive writing started with pen and paper (Pennebaker and Chung, 2007), but the practice has expanded to the information technology (Hardey, 2002). More recently, mobile apps (Wang et al., 2018; Wu and Bailey, 2018; MacIsaac et al., 2022) and chatbots (Howe et al., 2022; Park et al., 2021) have also been investigated to support expressive writing. For example, smartphone apps or chatbots ask users to type their negative emotions (Howe et al., 2022; Abd-Alrazaq et al., 2021; Ahmed et al., 2022). Smartphones, in particular, are now highly ubiquitous, which creates an opportunity to investigate and distribute affordable, effective, and easily accessible stress-lowering tools to those suffering from stress.
Due to recent technological advances, speech input is now a viable input option on smartphones. Previous HCI work has found that speech input has an advantage in speed (Ruan et al., 2018), workload (Rzepka et al., 2022), and content richness (Luo, 2022) over keyboard input. Further, studies about expressive writing without smartphones or chatbots have found that verbal expression had a more positive effect than written expression (e.g. (Esterling et al., 1994; Murray and Segal, 1994)). Speech-enabled expressive writing with smartphones is currently underexplored, as most existing expressive writing support systems rely on keyboard text entry (see e.g. (Howe et al., 2022; Park et al., 2021)). This warrants further studies into the suitability of speech input in expressive writing.
Another powerful aspect of smartphones is their high programmability, allowing developers to build experiences that mimic those outside the digital world. Clinical psychology has shown that deleting objects tied to emotion helps to resolve negative moods (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021). People may write a note with their negative emotion and put it into a shredder or simply tear it up by hand (Nussbaum, 1995; Soesilo et al., 2021). This consequent action is a widely recognized practice to obtain an explicit feeling of removing or destroying the negative experience (Briñol et al., 2013; Lee and Schwarz, 2011). Similar features can be implemented as digital mobile experiences through different visual feedback solutions, but the potential of these solutions has not well examined.
This paper aims to better understand the cathartic effect of expressive writing on mobile devices. We first examined the impact of speech input versus typed input on user experience and the cathartic effect within a mobile expressive writing application. We also adapted key findings from the first investigation and improved the mobile solution. We then compared different visual feedback designs that are based on prior work and were provided immediately after the users engaged in expressive writing. Specifically, these studies answer two research questions:
RQ1.
How do speech and text-based input in mobile expressive writing influence the stress-relieving effect, the content of expression, and the perceived usability and preference?
RQ2.
How do different visual feedback designs in mobile expressive writing influence stress-relieving effect, and the perceived usability and preference?
This work conducts two in-the-wild studies (15 and 23 days long with 28 and 36 participants, respectively) using a mobile chatbot application and offers the following contributions:
Investigation of the effects of input modalities and visual feedback designs provided after emotional disclosure in mobile expressive writing.
Design and research implications based on our two studies, highlighting future research directions in technology-supported expressive writing and stress coping.
2. Related Work
Our work is contextualized within the intersection of digital tools for expressive writing (Pennebaker, 1997; Pennebaker and Chung, 2007) and related HCI aspects.
2.1. Expressive Writing
The therapeutic use of writing is called expressive writing (Pennebaker, 1997; Pennebaker and Chung, 2007) (also called written emotional disclosure (Frisina et al., 2004) or therapeutic writing (Wright and Chung, 2001)). In expressive writing, people write freely about their traumatic or stressful events with their emotions in detail. Undergoing this expressive writing session is known for its various benefits in the long term, ranging from physical and mental health to social and behavioral aspects (Pennebaker, 1997; Reinhold et al., 2018). Expressive writing can bring increased distress and negative mood right after writing as short-term effects (Esterling et al., 1999) while prior research also found a decrease in depressive symptoms (Reinhold et al., 2018). To unfold these complex and contradictory outcomes, there has been a large discussion on the process underlying expressive writing.
The mechanism of how expressive writing works has not been fully understood, but there are three common theories to explain: inhibition theory, exposure theory, and cognitive processing (Reinhold et al., 2018). Early work on expressive writing followed inhibition theory, assuming that inhibiting thoughts or emotions caused a high mental workload, and therefore, expressing thoughts and emotions led to physical and mental health (Pennebaker and Beall, 1986). Another explanation, exposure theory, proposed that expressive writing might work as an imaginative exposure to negative emotional events and let people inured to stimuli. During exposure, the response to the stress was unlearned or replaced by other responses, and this process might relate to the outcome of expressive writing (Sloan and Marx, 2004). Cognitive processing spotlighted the content of expressive writing and claimed that the process of constructing a coherent story helped people organize their experiences and thoughts and change perspectives (Pennebaker and Chung, 2007). Considering these theories, expressive writing lets us confront and unleash our negative experiences and emotions, which we tend to inhibit or avoid. Although it may temporarily provoke negative moods, it still helps people restructure their cognition toward stressful events and brings various benefits.
Along with psychology revealing its effects and mechanisms, existing research has explored the applications of information technology to support expressive writing (Sauter, 2014; Hardey, 2002). Experiments with EmotionDiary showed how the tool was most effective for people with past experience on emotional disclosure on social networking services (Lee et al., 2016). Another group designed chatbot interactions for expressive writing, and the responsive chatbot enhanced emotional disclosure than writing on a document (Park et al., 2021). Other applications support journaling (Wang et al., 2018), reflection (McDuff et al., 2012) or recollection (Peesapati et al., 2010) of everyday emotional experiences as well as trauma and stress experiences.
One previous study explored chatbot designs to mitigate workplace stress by comparing three types of interventions: taking a break, supporting mindfulness, and emotional expression (Howe et al., 2022). While the emotional expression intervention reduced stress most efficiently, participants rated it the lowest because of two drawbacks: it required the highest effort, and they felt more stressed while reflecting their stress and the associated negative emotion. Technology-supported expressive writing, like conventional longhand writing, helped disclose emotions and reduce stress, while short-term deterioration of emotions was observed as well, and its burden became problematic. Therefore, there is a need to explore usability aspects of interface designs for mobile expressive writing for cathartic effect.
2.2. The Effect of Input Modalities in Expressive Writing
Previous work found that verbal expressions (e.g., speaking into a tape recorder) could improve cognitive change and enhance immune function more (Esterling et al., 1994) than written expressions (e.g., writing on a piece of paper) while both verbal and written expressions led to increase in negative emotion (Murray and Segal, 1994). For the degree of emotional expression, there are two conflicting results; one reported greater expression in written mode (Esterling et al., 1994) while the other in a verbal condition (Murray and Segal, 1994). Prior work has also examined the effects of verbal and written forms only in self-disclosure but also in disclosure to others. Iacovelli et al. (Iacovelli and Johnson, 2012) found that face-to-face disclosure showed stronger physiological improvements than remote communication via text messaging. Other work has compared therapy interviews (in verbal form with therapists) with expressive writing (in written form by oneself) and found that therapy interviews helped a shift to positive affect (Donnelly and Murray, 1991; Murray et al., 1989) and cognitive change (Murray et al., 1989). Combining these findings, verbal expression, which is the most natural and prevailing modality for therapy, may have a greater effect on post-expression mood and stress reduction than written expression.
Many of the existing expressive writing technologies and studies have mainly employed text entry using keyboards (either physical or virtual) (Abd-Alrazaq et al., 2019; Park et al., 2021; Howe et al., 2022). Brewin and Lennard’s study on comparing typing against writing for a long term revealed that typing resulted in less negative effects but less self-rated disclosure (Brewin and Lennard, 1999). This result motivates more extensive explorations on the effect of input modalities in technology-enabled expressive writing. Speech-based input is in particular a promising text entry approach commonly available in various smartphones and mobile devices (Enge, 2020). Speech-based input has been shown as superior over keyboard-based methods for entering short messages on touchscreen mobile devices (Ruan et al., 2018). It also has demonstrated benefits in information search with conversational agents (Rzepka et al., 2022) and food journaling (Luo, 2022). However, its effect on mobile expresstive writing is still underexplored. Comparative studies in input modalities thus contribute to deepening the understanding of the effect of technology-enabled expressive writing.
2.3. Feedback Effects in Expressive Writing
Prior work in psychology has shown that additional actions on the expressed negative emotion, such as physically throwing it away, can enhance the sense of stress relief (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021). These studies show that the physical clearing of materialized emotions leads to the psychological clearing of emotions. Washing or sanitizing hands can also diminish the extent and impact of previous negative sensations (Xu et al., 2012) or feelings of guilt (Zhong and Liljenquist, 2006). These psychological effects depend on the intensity of the physical action or the degree of embodiment. Soesilo et al’s experiment found that shredding the paper with scissors reduced envy more than tearing by hand (Soesilo et al., 2021). Briñol et al. (Briñol et al., 2013) also tested the effect of interaction on a computer, and revealed that moving files with negative emotions to the trash with a mouse had a greater effect than merely imagining them. As our cognition is based on sensory organs, the mental state is grounded in the physical state (Barsalou, 2008; Lee and Schwarz, 2021). Therefore, the embodiment and agency of the elimination action are important, but the effects when others perform the elimination movements are not fully clarified (Lee and Schwarz, 2021).
Digital technology offers great flexibility on the consequent actions on expresstive writing. Grieger et al. (Grieger et al., 2021) demonstrated that in virtual reality, punching objects displaying stressful messages promoted relaxation and encouraged self-reflection. Björling et al.(Björling et al., 2019) developed a virtual reality system to reduce stress in teenagers, which allowed users to burn their negative thoughts with leaves in a fire. While previous studies in virtual reality explored the effects of eliminating emotional objects, most studies and applications on mobile devices have implemented no feedback or only text feedback (Park et al., 2021; Howe et al., 2022). Yet, the effects of the visual feedback after emotional disclosure on mobile devices remain relatively underexplored.
3. Study 1: Examination of Input Modalities
We first conducted a study to explore the influence of using different input methods to answer RQ1. The following study protocol was approved by the Institutional Review Board of the first author’s institution.
3.1. System Implementation
We developed a chatbot system that allows users to perform expressive writing on their smartphones (Figure 1). We adopted chatbot interaction as it can enhance emotional disclosure rather than writing on a document (Park et al., 2021). We chose LINE as the underlying platform because it is one of the most widely used messaging apps in Japan, the country where we conducted our studies. In our implementation, participants could express their stressful experience that happened that day and how they felt then, using a given input method (Figure 1b). The instructions were based on the standard wording (Pennebaker and Chung, 2007), but translated into more casual, conversational words in reference to the study that designed a chatbot for expressive writing (Park et al., 2021). After the system received stress-related emotional expression content, it asked participants to rate the perceived stress changes at that moment (Figure 1e). The system could also send a message to remind participants and encourage them to share their experiences. We prepared fixed messages for notifications and responses. We prepared fixed notification and response messages for the chatbot, which also deliberately included emojis as prior work had shown the effect of making a chatbot more credible (Beattie et al., 2020) or enjoyable (Fadhil et al., 2018).
[Two screenshots of the Chatbot system interface we used in Study 1.]Two screenshots of the Chatbot system interface we used in Study 1. The first screenshot with areas labeled (a) to (c), and another screenshot with areas labeled (d) and (e). Section (a) dispalys the general chat UI of the system, showing the message sent from the system on the left side and message users sent on the right side. Section (b) displays the message sent by the system saying: ”Good evening. How was your day? Please tell us about a stressful event that happened today, and how you felt about it. No matter how trivial the event may seem, we ask you to describe that event using text input only. Please avoid using voice-note or other input methods. You may input your response by clicking on the keyboard symbol in the lower-left corner of the screen. We kindly ask that you submit your response by 24:00. Thank you for your cooperation.” Section (c) displays a menu at the bottom of the screen with two buttons: ”Evaluate your Stress Level” with a five-star icon, and ”Describe a Stressful Event” with a pen and paper icon. Below these buttons is text saying, ”* These buttons are only available if you have responded to our most recent survey.” and a clickable link saying, ”Click here to input your response”, followed by a ”Optional Menu” button. Section (d) shows a reply message to the expressive writing that says, ”Thank you for sharing this stressful event. We understand that some of this information may be difficult to share, but we greatly appreciate your cooperation. We kindly ask you to evaluate your current stress level once more.” Section (e) shows a message for ”Stress Evaluation” with buttons on it. The message starts with ”Comparing your responses before and after you described a stressful event, how do you feel your stress level has changed? Please press the button that corresponds with your response only one time.” Below this text are five buttons labeled ”Greatly Increased” in red, ”Increased” in light red, ”No Change” in gray, ”Decreased” in light green, and ”Greatly Decreased” in green. A note at the bottom reads, ”* As only responses received by 24:00 will be considered valid, we ask for your cooperation in responding as soon as possible. * Please make your selection only one time. We ask for your patience as our system may take some time to respond to you.”
3.2. Input Modality Conditions
We asked the participants to type text with the soft keyboard in keyboard-based condition and send a voice message in voice condition. In Japan, where this study was conducted, there are different keyboard designs people commonly use besides the QWERTY layout (e.g., flick input). We allowed our participants to use any keyboard they liked to maintain the realistic setting of their technology use. For the voice message input modality, we used the voice message functionality already available in LINE, with which our study participants audio-recorded their utterances. We prohibited our participants from using speech recognition functionalities for text input as the LINE platform did not offer a way to differentiate between text generated by speech recognition or text entered with a software keyboard.
3.3. Experiment Design
3.3.1. Independent Variables
We prepared two conditions for this part of the study: keyboard-based and voice conditions. We chose these two methods as they represent major text entry approaches commonly available in computers and smart devices (Yu et al., 2019). In using smartphones, soft keyboards are the most common text entry method while voice input is increasing its share (Enge, 2020). They also represent two common communication modalities in real-world therapy (i.e., writing down stressful experiences vs. talking to a therapist).
3.3.2. Dependent Variables for Stress-relieving Effect and Content
We captured the following four quantitative measures. For the text-related measures, we used Whisper111We decided to use Whisper because it exhibited a high level of recognition accuracy (approximately 94%) for Japanese at the time of our analysis. (Radford et al., 2022) to transcribe the collected voice data.
Immediate Stress Change
To capture how the participants perceived the cathartic effect of each expressive writing, we asked them to respond to the question of “Comparing your responses before and after you described a stressful event, how did you feel your stress level has changed?” (Figure 1e). This question was based on previous work that captured stress level change through stress coping interventions (Tong et al., 2023). The response was in a five-point Likert scale, ranging from [-2, 2] (2: greatly increased, 1: increased, 0: no difference, -1: decreased, and -2: greatly decreased). We asked participants to respond to this question immediately after every expressive writing.
Post-experiment Cathartic Effect
To capture the post-experiment perceptions of the cathartic effect, we asked “{Keyboard-based/Voice} input was helpful for reducing my stress,” based on Howe’s work (Howe et al., 2022). The response was a five-point Likert scale (1: strongly disagree to 5: strongly agree). We collected this data in the post-experiment questionnaire.
Length
We counted the length of each expressive writing. The length was calculated as the character length of the original Japanese text, and one English word experimentally corresponds to 2–2.5 characters in Japanese in terms of word/character counts.
Emotional Words
To capture how much the participants disclosed their emotions in the expressive writing, we collected the use of words related to negative emotion. The usage of emotional words is known to relate to cathartic effect (Pennebaker and Chung, 2007). We analyzed the text with the Linguistic Inquiry and Word Count program (LIWC) (Pennebaker et al., 2001; Igarashi et al., 2022). For each participant/input modality combination, we separated the descriptions into words using MeCab (Kudo, 2005) (a widely used segmentation and morphological analysis tool for Japanese) and then fed them into LIWC. We used the score of positive/negative emotion words from the results of LIWC analysis, which is the percentage of words registered in the library as positive/negative emotion words out of the number of words that appeared in the text.
3.3.3. Dependent Variables for Usability and Preference
We collected the following three quantitative data measures in the post-experiment questionnaire to understand how the participants perceived the usability of each input modality and which modality they preferred.
Ease of Use
To capture perceptions of the ease of use, we asked a question of “It was easy to enter stress by {Keyboard/Voice},” based on previous work compared the usability of digital coping methods (Howe et al., 2022). The response was a five-point Likert scale (1: “It was difficult to do” to 5: “It was easy to do.”).
Mental Workload
To capture the perceived mental workload associated with each input modality condition, we used NASA-TLX (Hart and Staveland, 1988). NASA-TLX’s score ranges from 0 to 100, calculated as the weighted sum of the six subjective subscales.
Preference
To capture the preference for input modalities, we asked a question of “Which input method would you prefer to use?” with two answering options: “keyboard-based input” and “voice input.”
3.4. Procedure
On the first day, we asked the participants to sign the consent form, fill in a pre-experiment questionnaire (see Appendix A), and ensure that they were able to use our system. The pre-experiment questionnaire included the Perceived Stress Scale (PSS) (Cohen et al., 1994) to gauge the stress levels of our participants prior to the experiment. We explained that the main objective of this study was to collect their stress-related emotion through our system instead of investigating the effect of the two input modalities. In this manner, we avoided giving participants potential biases. After the pre-experiment questionnaire, we shared the link and QR code for the installation of our chatbot. We split participants into two groups with considering the balance of gender, age, and occupations to counter-balance the order of the presentations of the two input modality conditions.
From Day 2 to Day 14, our system asked participants to share their stressful experiences and related emotions given one of the input modalities. We sent the question between 6 p.m. and 7 p.m. daily, and reminded the incomplete participants between 9 p.m. and 10 p.m. Additionally, since stress has been noted to have correlations with weekdays (Brantley et al., 1988), we aimed to minimize this effect by having participants experience two different conditions on a weekly basis. On Day 8, we flipped the input modality conditions. We set the condition lengths to 7 days and 6 days so that the first and final day, which required a relatively longer time to answer the questionnaire, occurred on weekends and made participation easier.
On the last day, we conducted a debriefing where we disclosed the true objectives of this study. We explicitly explained that they would be fully compensated even if they decided to withdraw at this point, and explicitly offered an opportunity to withdraw. None of the participants withdrew from the study. We then asked the participants to fill out a post-experiment questionnaire (Table 3 in Appendix A), which included questions about ease of use, mental workload and post-experiment cathartic effect of each input modality and the preference on input modalities, and the questions for PSS. We also collected qualitative data with the open-ended questions, asking participants for their thoughts or opinions about each input condition and their reasons for their preferences. We offered all participants approximately 19 USD in the local currency for the completion of the study.
3.5. Participant recruitment
We recruited participants through Crowdworks, one of the most widely used crowdsourcing platforms in Japan. Majima et al. (Majima et al., 2017) confirmed that the subjects available through Crowdworks are sufficiently reliable for behavioral research. We recruited 28 participants (14 female and 14 male; PA1 – PA28) in their 20s to 60s (Table 1).
| Groups | Gender | Age Ranges |
| Group 1 (Keyboard Voice) | F: 7, M: 7 | 20s: 2, 30s: 4, 40s: 3, 50s: 3, 60s: 2 |
| Group 2 (Voice Keyboard) | F: 7, M: 7 | 20s: 2, 30s: 4, 40s: 4, 50s: 3, 60s: 1 |
3.6. Data Analysis
To compare quantitative data collected in each condition or at pre- and post-experiment, we used non-parametric statistical tests. For analyzing qualitative data, two of the authors performed a thematic analysis (Braun and Clarke, 2006). The first author read through the free-text responses from the post-experiment questionnaire and developed a codebook with respect to factors that determined participants’ preferences. Then, the first author and another author reviewed the codes and discussed until a consensus was reached. Finally, we categorized the codes into themes.
4. Study 1 Results
We collected 483 expressive writing data and 479 data on immediate stress change, which recorded a response rate of 98.9%. Among 483 expressive writing data, 244 and 239 were entered under the keyboard text entry and voice message conditions, respectively. The average scores of pre- and post-experiment PSS were 28.5 ( = 7.87) and 27.6 ( = 8.95), respectively, and our Wilcoxon signed-rank test did not find a significant difference ( = 1, = 0.250, = .27, = 0.0473).
4.1. Perceived Usability & Preference
Ease of Use
The average values of ease of use were 4.11 ( = 1.10) and 2.46 ( = 1.57) for the keyboard text entry and voice message conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -4.18, ¡ .001, = 0.79).
Mental Workload
The average values of mental workload were 51.7 ( = 16.5) and 62.0 ( = 19.0) for the keyboard input and voice message conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -3.46, ¡ .001, = 0.65). There were significant negative correlations between the ease of use and the mental workload scores in both the keyboard text entry ( = -0.39, ¡ .05) and voice message ( = -0.43, ¡ .05) conditions.
Preference
Twenty-one participants preferred keyboard text entry to voice messages while seven preferred voice messages (Figure 2). Our chi-square test revealed that participants preferred the keyboard text entry significantly more compared to the voice message condition ( = 7.0, ¡ .01).
We further examined comments we obtained from our participants in the post-experiment questionnaire. We identified three major factors that highlighted the two important factors in the keyboard text entry method: privacy friendliness, and reflection through typing.
Our participants expressed that the voice message input was often inappropriate, because emotional disclosure tends to include private or sensitive information. Stressful events often occur on the job or in relationships with others, and expressive writing may include negative emotions toward others. Therefore, our participants hesitated to record voice messages in public spaces as PA1 mentioned: “I found it very challenging to use voice input because I didn’t want others to hear me, so I had to wait for moments when there were no people around. This was especially difficult when I was at work in the company.”
On the contrary, typing allowed participants to refine their thoughts and feelings because they could review and modify the content. As they typed on the keyboard, the text they entered appeared on their screen. They reviewed the content, making revisions as needed, which helped them review their writing objectively and organize their thoughts and emotions as PA12 commented “I was able to write about the events of the day while recalling them in detail, which allowed me to organize my emotions.”
In contrast, some participants said organizing the content while speaking was challenging, and they sometimes found themselves not knowing what to say in the middle of recording. PA1 commented: “I felt that I didn’t know what to say once I started voice input, and I couldn’t properly convey my thoughts.” With the keyboards, our participants were relaxed and could enter calmly: “Also, when it came to speech, I was too nervous to say what I was thinking, so texting was easier for me since I could slowly look back and express my thoughts.” (PA2).
4.2. Stress-relieving Effect & Content of Expression
Immediate Stress Change
The average values of immediate stress change were -0.14 ( = 0.44) and -0.085 ( = 0.50) under the keyboard input and voice message conditions, respectively (Figure 3). We ran a Mann-Whitney’s U test to evaluate the difference in the cathartic effect, but we could not find a significant effect ( = 27087, = -0.67, = .47, = 0.03).
Post-experiment Cathartic Effect
Figure 3 shows the average values of the responses to the question asking the cathartic effect (the keyboard text entry: 3.21, = 0.99; the voice message: 2.61, = 1.10). The keyboard text entry condition was rated significantly higher than the voice message condition, as indicated by our Wilcoxon signed-rank test ( = 1, = -3.85, ¡ .05, = 0.73). We also found that the post-experiment cathartic effect had a significant positive correlation with the perceived ease of use in both the keyboard text entry ( = .52, ¡ .01) and voice message conditions ( = .71, ¡ .001).
Length
The average length of emotional expression data was 65.2 characters ( = 43.8) under the keyboard condition and 63.6 ( = 38.4) under the voice message condition in Japanese 222Each data would correspond to roughly 25–32 words in English on average, for one English word experimentally corresponds to 2–2.5 characters in Japanese in terms of word/character counts.. Our Wilcoxon signed-rank test did not reveal a significant difference ( = 1, = 0.73, = 0.48, = 0.14).
Emotional Words
The use of words related to emotions also had no significant difference between the two conditions. Our Wilcoxon signed-rank test did not find a significant effect of input modality on the appearance frequency of positive nor negative emotion words (positive emotion: = 1, = -1.50, = .50, = 0.29; negative emotion: = 1, = -0.096, = .93, = 0.018). Input modality did not affect the content of expressive writing texts.
These four quantitative metrics led us a conclusion that the two input modalities did not exhibit clear differences with respect to stress-relieving effect as well as the content in expressive writing.
4.3. Summary
Our first study confirmed that the keyboard-based text entry provided better usability and was preferred more because of its privacy friendliness and reviewing process. On the other hand, the two input modalities did not show clear differences in immediate stress relief, and pre- and post-experiment stress levels showed no significant difference. We, therefore, kept working on improving the interface design for mobile expressive writing. We moved our attention from input to output and examined how visual feedback could impact perceived cathartic effects.
5. Study 2: Examination of Effects of Visual Feedback
Our next exploration is to examine the visual feedback provided after expressive writing, related to RQ2. We obtained a separate approval for the study protocol described in this section form the Institutional Review Board of the first author’s institution.
5.1. Experiment Design
We used the same system architecture with additions of visual feedback for the expressed emotional writing. We decided to use the a keyboard-based text entry method based on our findings in the first study. We also made a minor modification to allow participants to share positive episodes when they did not have any stressful moments though we did not use such data for our later analysis.
5.1.1. Independent Variables: Visual Feedback Conditions
We added two visual feedback conditions: Passive and Interactive. The two conditions differed by the extent of embodiment (Soesilo et al., 2021) and the agency of elimination (Lee and Schwarz, 2021). These two factors might influence the outcome, referring to the previous work (Lee and Schwarz, 2021). We also included the None condition where no visual feedback was provided as a baseline, which is equivalent to the configuration used in Study 1.
Passive Feedback Condition
This condition offered participants unembodied, non-interactive visual feedback. When participants submitted their stress-related emotional expression text, the system provided a short video where a note with the submitted text exploded. As we introduced in section 2.3, materializing negative emotions and giving feedback that the object extinguished also dissolve its tied emotions (Briñol et al., 2013). In the context of mobile stress management, users may rather prefer low-burden interactions (Howe et al., 2022), and physically-demanding actions may not be suitable. This condition also represents a situation where users delegate deleting actions to others, which has not been fully explored yet (Lee and Schwarz, 2021). Figure 4 shows a walkthrough of this condition. We adopted an explosion visual effect similar to feedback used in the previous VR studies (Grieger et al., 2021; Björling et al., 2019).
Interactive Feedback Condition
This condition offers participants embodied, interactive visual feedback. After submitting their stress-related emotional expression text, the system provides an interface where they can proactively explode a note with the submitted text by tapping it. This design was also based on studies on destroying materialized emotions. Figure 5 shows the walkthrough of this condition. This condition thus would require additional actions from users, but may help them build the sense of control, potentially contributing to cathartic effect positively.
5.1.2. Dependent Variables for Stress-relieving Effect
Similarly to our first study, we collected immediate stress change and post-experiment cathartic effect. But we decided not to collect the data related to the content of the expression as our focus was the feedback after the expression. We also obtained the following quantitative data for our analysis.
Long-term Stress Change
To gauge how the participants’ stress levels changed over a longer period, we used Perceived Stress Scale (PSS) (Cohen et al., 1994). This assessment tool consists of 14 questions about how often each certain situation occurred in the past week on a five-point Likert scale ranging from (0: never – 4: very often). After collecting the score every week, we subtracted the after score from the before score and used it as the long-term stress change for the condition they experienced that week.
5.1.3. Dependent Variables for Usability and Preference
We collected ease of use, mental workload and preference in this study as well. In addition, we collected the following effort measurement. All four were collected from the post-experiment questionnaire.
Effort
To capture perceptions of the effort required, we asked “This feedback requires effort,” based on previous work compared the usability of digital coping methods (Howe et al., 2022). The response was in a five-point Likert scale (1: “Not at all” – 5:“Strongly agree”).
5.2. Procedure
The study lasted for 23 days (three full weeks for data collection and two days for the onboarding and post-experimental stages) to accommodate the three conditions (None, Passive, and Interactive).
The process in the onboarding and post-experimental sessions was the same as in Study 1. In this study, we explained our participants the objective of our study to examine the effects of three feedback designs at this time. The pre-experiment questionnaire included the questions from PSS for long-term stress change. The onboarding required approximately 30 minutes to complete on the first day. We split participants into six groups with considering the balance of gender, age, and occupations to fully counter-balance the order of the presentations of the three feedback conditions.
Participants are under each of the visual feedback conditions for each week of the experiment (Day 2–8, Day 9–15, and Day 16–22). On the mornings of days 9 and 16, we asked participants to fill out PSS to track their long-term stress level trends.
We conducted a post-experiment questionnaire once on the final day. We then asked the participants to fill out a post-experiment questionnaire (Table 4 in Appendix B), which included questions for PSS and about post-experiment cathartic effect, effort, ease of use, and mental workload of each feedback condition and the preference on feedback conditions. We also collected qualitative data with open-ended questions, asking participants for their thoughts or opinions about the positive and negative aspects of each feedback condition and their reasons for their preferences. We also asked when they would like to use each condition and what other feedback designs they would like to have. The post-experiment questionnaires took around 30 minutes to complete on the final day. We offered all participants approximately 21 USD in the local currency for the completion of the study.
We performed data analysis in the same way as we did in our first study.
5.3. Participants
We used the same crowdsourcing service as Study 1. We also followed the same procedure to choose participants. As a result, we recruited 36 participants (18 women, 18 men; PB1 – 36) in their 20s to 60s (Table 2). All participants self-claimed that they did not know of the existence of our first study.
| Groups | Gender | Age Ranges |
| Group 1 (None Passive Interactive) | F: 3, M: 3 | 20s: 0, 30s: 3, 40s: 2, 50s: 1, 60s: 0 |
| Group 2 (None Interactive Passive) | F: 3, M: 3 | 20s: 1, 30s: 2, 40s: 2, 50s: 1, 60s: 0 |
| Group 3 (Passive None Interactive) | F: 3, M: 3 | 20s: 0, 30s: 3, 40s: 2, 50s: 1, 60s: 0 |
| Group 4 (Passive Interactive None) | F: 3, M: 3 | 20s: 1, 30s: 2, 40s: 2, 50s: 1, 60s: 0 |
| Group 5 (Interactive None Passive) | F: 3, M: 3 | 20s: 1, 30s: 2, 40s: 2, 50s: 0, 60s: 1 |
| Group 6 (Interactive Passive None) | F: 3, M: 3 | 20s: 1, 30s: 2, 40s: 2, 50s: 1, 60s: 0 |
6. Study 2 Results
We collected 550 expressive writing data with immediate stress change data, and the response rate was 99.4%. Among 550 expressive writing data, 186, 183, and 181 were under the None, Passive, and Interactive conditions, respectively.
6.1. Quantitative Results for Stress-relieving Effect
Long-term Stress Change
The average difference between pre- and post-experimental PSS scores were 0.39 ( = 5.23), -0.58 ( = 5.03), and -2.22 ( = 7.06) in the None, Passive, Interactive conditions, respectively (Figure 6). Our Friedman test did not reveal any significant difference ( = 1.22, = .54, = .117). For the Interactive condition, our Wilcoxon signed-rank test found a significant difference between pre- and post-experimental phases ( = 1, = 1.05, ¡ .05, = .175). As for the None and Passive conditions, our Wilcoxon signed-rank test did not reveal a significant difference between before and after (None: = 1, = 0.267, = .49, = 0.0445; Passive: = 1, = 0.0393, = .25, = 0.00655).
The average PSS scores in the pre- and post-experimental phases were 24.75 ( = 7.93) and 22.33 ( = 8.66), respectively (Figure 6). Our Wilcoxon signed-rank test revealed a significant difference the pre- and post-experimental phases ( = 1, = 0.448, ¡ .05, = 0.0746). Experiencing mobile expressive writing through the system for three weeks thus significantly reduced the stress.
Immediate Stress Change
The average scores of immediate stress change were -0.40 ( = 0.46), -0.49 ( = 0.55), and -0.57 ( = 0.48) under the None, Passive, and Interactive conditions, respectively (Figure 6). Our Friedman test did not reveal any significant difference ( = 2.64, = .27, = 0.25).
Post-experiment Cathartic Effect
We also examined the ratings of post-experiment cathartic effect. Figure 6 shows the average ratings of all the three conditions: 3.44 ( = 0.97), 3.75 ( = 1.11), and 3.5 ( = 1.13) under the None, Passive, and Interactive conditions respectively. Our Friedman test did not reveal a significant difference ( = 2.14, = .34, = 0.21).
6.2. Qualitative Results for Stress-relieving Effect
6.2.1. Two Visual Feedback Designed Provided a Sense of Catharsis but in Different Ways.
Our participants mentioned the cathartic effects as the advantages of Passive and Interactive conditions though our analysis found differences between them. The major difference emerged from the difference in agency of elimination. In the Interactive condition, the participants gained the cathartic effect as they exploded the letters by themselves. Seven participants felt like they were relieving their stress on their own: “I myself had exploded the stress, and it felt refreshing.” (PB32). The presence of interactivity was influencing agency in stress relief, and our participants pointed out the lack of agency as the disadvantage of the Passive condition. They felt they were “just like observing” (PB28) and not having control. In the visual feedback of stress disappearance, giving control over the vanishing provided participants with a sense of active control over the process, and it enhanced the cathartic effect.
Our participants saw the cathartic effects of the Passive condition by gaining an opportunity to take a distance from the experienced stress. Five participants mentioned laughter as the advantage of the Passive condition: “It was good that it was simply amusing and made me laugh. I thought it was especially good when I was feeling stressed out and bothered.” (PB11) This explicit visualization of the explosion of the externalized stress-related emotion also made the participants feel refreshed or calm down: “Seeing the stress words explode and disappear made me feel a bit amusing in a ‘deserve it!’ kind of way, and I appreciated that it made me feel calm.” (PB30).
6.2.2. No Feedback Still Offered Reflection Opportunities.
Our qualitative results did not find explicit benefits of catharic effects in the None condition. PB30, for example, pointed out its disadvantage of not offering an explicit way to vent: “When there is too much stress, thinking about which ones to list or write all of them, or remembering that stress, causes extra stress. I feel like it’s a waste of time to dwell on it, remember it, and go back and forth over it. That’s what I find problematic.”
However, our participants saw the value of this condition outside the cathartic effect. Participants commented that the system was actively listening to their expressions in this condition. They appreciated the immediate replies from our system, which may relate to this effect. The reply message333The message used in the second user study was “Thank you for sharing this stressful event. We understand that some of this information may be difficult to share, but we greatly appreciate your cooperation.” was pre-defined and used across all the conditions. Some of our participants considered this reply as feedback to them as it offered them a sense of confirmation that they were being heard: “It was beneficial to be able to directly describe the accumulated stress of the day to the chatbot. The immediate response from the chatbot made me feel like it was actively listening to my thoughts.” (PB26)
Organizing Thoughts
Another benefit perceived by our participant was to help participants organize their thoughts. Seven participants mentioned that they could organize their thoughts and objectively recognize and contemplate their stress: “I was able to reflect on the events that caused stress and think about them logically. It was beneficial to calmly reconsider the origins of feelings of aversion and anger towards others. It was beneficial to calmly reconsider the origins of feelings of aversion and anger towards others.” (PB11) This result suggests that the absence of the visual feedback may encourage reflection of participants’ negative feeling, promoting another benefit of expressive writing.
6.3. Quantitative Results for Usability & Preference
Ease of Use
The average ratings of ease of use were 4.06 ( = 0.83), 3.92 ( = 0.91), and 3.78 ( = 1.10) under the None, Passive, and Interactive conditions respectively (Figure 7). Our Friedman test did not show a significant difference ( = 0.42, = .81, = 0.04).
Effort
The average ratings of effort were 2.36 ( = 1.02), 2.00 ( = 1.15), and 2.56 ( = 1.18) under the None, Passive, and Interactive conditions respectively (Figure 7), confirming a significant difference with our Friedman test ( = 10.18, ¡ .01, = 0.98). The post-hoc Nemenyi test confirmed significant differences between Passive and Interactive ( ¡ .05).
Mental Workload
The average mental workload scores were 43.4 ( = 19.5), 39.0 ( = 19.67), and 46.35 ( = 17.95) under the None, Passive, and Interactive conditions respectively (Figure 7). Our Friedman test revealed a significant difference ( = 7.90, ¡ .05, = 0.760). The post-hoc Nemenyi test confirmed significant differences between Passive and Interactive ( ¡ .05).
Preference
In the post-experiment questionnaire, 9, 14, and 13 participants mentioned the None, Passive, and Interactive conditions as the most preferred feedback, respectively. Our chi-square test did not reveal that participants preferred specific conditions ( = 1.17, = 0.56).
6.4. Qualitative Results for Preferences
6.4.1. Scenario-dependent Use of Feedback Designs
Our participants shared their views on different use scenarios for each design. They commented that the None condition could be useful when they simply wanted someone else to hear their complaints. Another notable use scenario was for reflecting on their feelings through writing or leaving records for future reflections: “I want to use it (the None condition) when I can’t talk to others about something, but it’s hard for me to keep it to myself, and I want to be heard, even if it’s by a robot.” (PB29)
With respect to the Passive condition, our participants expressed that they would use it when the stress originated from others. Other participants also explicitly mentioned that they would rather avoid using the Passive design when the cause of their stress was associated with themselves: “I want to use it (the Passive condition) when I feel stressed about others. I would not use it for stress within myself because I don’t feel it is very effective.” (PB13)
The Interactive condition was deemed useful when participants felt intense stress. However, our participants felt that they would not use it when they were exhausted: “I wanted to use the Interactive condition when I was feeling a bit strongly stressed. There are quite a few times when expressions alone are not enough to completely relieve stress, so I felt that this condition would help me relieve even the intense stress.” (PB33) This impression is in line with our quantitative results on the perceived effort, where the Interactive condition was rated as significantly more effortful than the Passive condition (Section 6.3).
In summary, our qualitative results highlight different possible use scenarios of the three visual conditions that the quantitative results did not uncover. Our participants considered that the None condition would be rather useful for seeking mental cool-down. Their comments also clarified the similarities and differences between the Passive and Interactive conditions. Both feedback designs were considered better suited for stress stemming from others, while the Interactive design can be effective for supporting the relief of intense stress.
6.4.2. Further Explorations of Visual Feedback Designs
Our participants offered various ideas on other possible visual feedback designs. They are categorized into three design philosophies: designs that offer immediate destruction; designs that offer gradual disappearance; and designs that offer detachment. Designs to show immediate destruction included braking blocks and letting virtual characters eat.
Designs to show gradual disappearance are visual effects where expressed negative experience would slowly disappear, such as a fade-away effect of the written text. Our participants commented that such effect may offer more calm feeling for stress relief: “I felt a sense of calm with a video that had a fading text effect, and I would like to continue using it. I thought the explosion effect was too exciting for stress relief and not very effective.” (PB30)
Designs to show detachment intend to allow users to simply leave their negative experience outside themselves. This design philosophy matches with the None condition, which may promote reflection. Our participants suggested a calm visual feedback that is not directly related to the written experience, such as flowers blooming up and animals playing around: “I think something soothing or relaxing, like animals, would be good.” (PB8)
7. Discussion
7.1. RQ1: Advantages of Keyboard-based Text Entry for Stress-related Emotional Expression
Consistent with previous research in psychological literature (Murray and Segal, 1994), our results from the first study confirmed that both speech and text input produce similar immediate cathartic effects. This result helps to understand the advantage of therapy interviews to expressive writing (Donnelly and Murray, 1991; Murray et al., 1989), which may lie in not the mode of expression (speech vs. written) but the presence of the listener.
Our results also showed that Keyboard-based text entry had an advantage in usability and preference when compared to Voice message input. This contradicts the previous studies (Kim et al., 2018; Luo, 2022; Rzepka et al., 2022; Kocielnik et al., 2018) indicating the superiority of speech input to keyboard input in usability and preference in other use contexts. Our qualitative results suggest two possible explanations: the confidentiality and the affordances of the keyboard-based entry are a better fit for expressive writing that can involve private matters.
Stress-related expression can contain highly-private content. In recording voice messages, our participants spoke their thoughts aloud in various settings, including in public spaces. This privacy threat has been a major concern of voice assistants and depends on the type of information and the location (Easwara Moorthy and Vu, 2015). However, these concerns may not be particularly serious in case where users utter non-private content (e.g., commands (Kim et al., 2018) or short journaling (Luo, 2022)) in private situations. As discussed in Section 4.1, our participant recognized the value of keyboard-based input even in private settings as it allowed participants to relax, write at their own pace, and avoid additional strain or stress. Our findings thus suggest that effects by input modality differences can form important application-dependent effect, stimulating further application-focused research
Our results also suggested that keyboard-based text entry may have helped the cognitive restructuring with its affordances. The keyboard-based condition was preferred because of its persistence and editability. These two affordances (Treem and Leonardi, 2013) may have enhanced the effective process of expressive writing: exposure and cognitive processing. Typing can make the explanation of one’s stress visible and reviewable persistently to oneself (Clark and Brennan, 1991). This may have worked as exposure to their past stressful events, a significant process in expressive writing (Sloan and Marx, 2004). Editability offers a capability to recraft and establish an ideal explanation (Walther, 1993). Editing the text during expressive writing may encourage people to reconsider our thoughts and reconstruct perceptions, known as cognitive processing, another important process to change our perspectives positively (Pennebaker and Chung, 2007). Keyboard-based text entry was thus able to successfully offer desirable features for mobile expressive writing.
7.2. RQ2: Visual Design Preferences Affected by Perceived Cathartic Effects, Sense of Agency, Required Effort, and Sense of Self-Denial
Our results in the second study unveiled different benefits and possible use scenarios of the three visual feedback conditions. This result is not fully aligned with prior work showing the advantages of performing post-writing tearing that create a feeling of destruction (Briñol et al., 2013; Grieger et al., 2021). One possible explanation for this contradiction is that stress relates to diverse conditions and emotions. According to our results, the cause of stress, particularly whether it comes from users themselves or others, was one of the determining factors to the perceived appropriateness of feedback designs. Previous studies only examined emotion about specific topics, such as one’s body or Mediterranean diet (Briñol et al., 2013) or envious experience (Soesilo et al., 2021). Previous literature also reported the influence by the source of emotion. Briñol et al. found that negative emotions related to participants’ bodies persisted even after tearing up their written thoughts, while negative emotions about the Mediterranean diet disappeared (Briñol et al., 2013). Grief therapy research has examined the relationship between the type of emotion and the way of disposal (Sas et al., 2016). Dynamic disposal (e.g., burning) is suitable for anger, but open or covert disposal (e.g., drifting or burying) is rather appropriate for sadness or regret. Our results suggest that future systems should combine different feedback designs to accommodate people’s various types of stress and emotions, offering an important direction for future research.
Participants also found that the None design was still useful for calmly reflecting on their emotions. This might be another explanation for our quantitative results not showing differences in the perceived cathartic effect. The reply to the expressive writing may have functioned as active listening. Active listening is one of the key skills in counseling (Hutchby, 2005) and helps speakers be more emotionally mature and open to their experiences (Rogers and Farson, 1957). Our results suggest that future research on stress management systems should consider purposeful absence of visual feedback to encourage users’ focus on their narrative and promote cognitive restructuring.
7.3. Research Opportunities in Mobile Expressive Writing
The affordances of modern smartphones are different from those of pen and paper — the original tools of expressive writing. Prior work has shown that expressive writing helps release repressed emotions, reorganize feelings and calm down (Pennebaker and Chung, 2007). Yet, most of past work has naturally been conducted with tools other than the relatively modern ones, such as smartphones. Our results with digital smartphone-powered expressive writing thus contribute to new empirical understanding of how interface factors of mobile technology can influence expressive writing effect and experience.
The qualitative findings in the second study suggest that different visual designs offer different experiences for the users. People felt a different level of agency over relieving their stress; participants felt as if they are observers in the Passive condition while they felt like they are actively destroying their stress in the Interactive condition. Existing literature suggests that personality (e.g., extrovert vs. introvert) could influence preferences and use scenarios of different visual feedback (Carver and Connor-Smith, 2010). A more extensive study considering these and other similar factors will lead into a deeper understanding of appropriate interface designs for different users and use cases.
The latest speech-based mobile input technologies can perform real-time recognition and transcription (Li et al., 2022). Engaging in a sensitive task, such as expressive writing, calls for exploring optimal input methods. Our results between text and speech based input showed clear differences, as explored in Section 4.1. Related to this, Luo demonstrated a combination of keyboard-based and speech-based text entry for food journaling (Luo, 2022), and such mixture may obtain a more optimal experience with respect to usability and stimulation of self-reflection. Examining the division of input tasks between keyboard-based and speech-based text entry for mobile expressive writing is yet another an interesting future research direction.
Overall, our results suggest that smartphones can facilitate effective expressive writing. Given their unique features, they can facilitate many types of rich interactions that are not available with other writing platforms. Auditory or haptic feedback (Takami et al., 2022) may be a possible direction to explore with expressive writing solutions For instance, Costa et al. (Costa et al., 2016) showed how haptic feedback of heart rate could help users regulate their anxiety. Future systems could include deep breathing and mindfulness activities with emotional expression and examine how they could enhance cathartic effects and overall stress relief in conjunction with expressive writing tasks. Such multimodal, multi-approach experiences are fairly trivial to program with modern smartphones coupled with off-the-shelf wearable sensors, such as heart rate monitors. Multimodal interface and feedback designs for stress relief may thus lead to a new interdisciplinary research opportunity for HCI and psychology.
7.4. Limitations
We acknowledge limitations in our work. We conducted our studies in Japan, and they required proficient Japanese communication skills. Accordingly, all participants were Japanese, constituting a homogeneous sample pool with respect to cultural backgrounds. This cultural specificity may have influenced the effectiveness of speech input as talking does not strongly relate to thinking in some Asian cultures, unlike in Western contexts (Kim, 2002). Future studies should examine how speech input could offer different effects and user experience in other cultural contexts. This work would serve as a reference for such studies to uncover the influence of cultural differences on mobile expressive writing.
Another homogeneity in our participants was that none of our participants regularly used speech input with their smartphones, which might have affected the preference results. However, our qualitative results uncover reasons of preferences that go beyond the established familiarity to software keyboards, and thus confirm the advantage of keyboard-based text entry for mobile expressive writing. Our results encourage researchers to further explore input modality effects in different cultural backgrounds and user populations.
The studies reported in this article are also rather short, compared to some of the seminal work in expressive writing in the field of psychology. Yet, during the time span our participants used the developed solutions, we were able to collect sufficiently data for quantitative and qualitative account of how the input modalities and visual designs differed. We argue these studies adequately contribute to the HCI literature on exploring smartphones as an expressive writing tool for cathartic effect, even if not so much on the already well-established method of expressive writing itself.
8. Conclusion
Expressing negative emotions is known to be an effective stress-coping mechanism, but challenges with respect to interface designs emerge when the practice is transitioned to technology-supported mental well-being systems. This work investigated how input modalities for emotional expression and visual feedback designs for post-expression interventions could affect the perceived cathartic effects and user experience. Our study highlighted the benefits and stronger preference for keyboard-based text entry over input methods using voice messages, and also revealed how different visual feedback designs can benefit users to accommodate different scenarios of stress-related emotional expression. Our findings also suggest important design considerations and future research directions in mobile expressive writing, and further encourage interdisciplinary research between HCI and self-help approach in psychology.
Acknowledgements.
We are grateful for the support received during the preparation of this paper. In particular, we would like to thank Ginshi Shimojima, Chi-lan Yang, Zefan Sramek, and Shitao Fang for their valuable feedback. We also extend our appreciation to the participants in our user study for their involvement. This research is part of the results of Value Exchange Engineering, a joint research project between R4D, Mercari, Inc., and the RIISE. This project was supported by JSPS Bilateral (Grant Number: 120232701).References
- (1)
- Abd-Alrazaq et al. (2019) Alaa A Abd-Alrazaq, Mohannad Alajlani, Ali Abdallah Alalwan, Bridgette M Bewick, Peter Gardner, and Mowafa Househ. 2019. An overview of the features of chatbots in mental health: A scoping review. International Journal of Medical Informatics 132 (2019), 103978.
- Abd-Alrazaq et al. (2021) Alaa A Abd-Alrazaq, Mohannad Alajlani, Nashva Ali, Kerstin Denecke, Bridgette M Bewick, and Mowafa Househ. 2021. Perceptions and opinions of patients about mental health chatbots: scoping review. Journal of medical Internet research 23, 1 (2021), e17828.
- Ahmed et al. (2022) Arfan Ahmed, Sarah Aziz, Mohamed Khalifa, Uzair Shah, Asma Hassan, Alaa Abd-Alrazaq, and Mowafa Househ. 2022. Thematic analysis on user reviews for depression and anxiety chatbot apps: Machine learning approach. JMIR Formative Research 6, 3 (2022), e27654.
- Barsalou (2008) Lawrence W Barsalou. 2008. Grounded cognition. Annu. Rev. Psychol. 59, 1 (2008), 617–645.
- Beattie et al. (2020) Austin Beattie, Autumn P Edwards, and Chad Edwards. 2020. A bot and a smile: Interpersonal impressions of chatbots and humans using emoji in computer-mediated communication. Communication Studies 71, 3 (2020), 409–427.
- Björling et al. (2019) Elin A Björling, Rachael Cicero, Aditya Sankar, and Anand Sekar. 2019. Thought Disposal: Co-Designing a virtual interaction to reduce stress in teens. In Proceedings of the 18th ACM International Conference on Interaction Design and Children. 562–567.
- Brantley et al. (1988) Phillip J Brantley, Thomas B Cocke, Glenn N Jones, and Anthony J Goreczny. 1988. The Daily Stress Inventory: Validity and effect of repeated administration. Journal of Psychopathology and Behavioral Assessment 10 (1988), 75–81.
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Breuer and Freud (1955) Josef Breuer and Sigmund Freud. 1955. The Standard Edition of the Complete Psychological Works of Sigmund Freud. Translated from the German Under the General Editorship of James Strachey. In Collaboration with Anna Freud. Assisted by Alix Strachey and Alan Tyson. Vol. 2. 1893-1895. Studies on Hysteria, by Josef Breuer and Sigmund Freud. Hogarth Press.
- Brewin and Lennard (1999) Chris R Brewin and Hayley Lennard. 1999. Effects of mode of writing on emotional narratives. Journal of Traumatic Stress: Official Publication of The International Society for Traumatic Stress Studies 12, 2 (1999), 355–361.
- Briñol et al. (2013) Pablo Briñol, Margarita Gascó, Richard E Petty, and Javier Horcajo. 2013. Treating thoughts as material objects can increase or decrease their impact on evaluation. Psychological science 24, 1 (2013), 41–47.
- Carver and Connor-Smith (2010) Charles S Carver and Jennifer Connor-Smith. 2010. Personality and coping. Annual review of psychology 61 (2010), 679–704.
- Clark and Brennan (1991) H. H. Clark and S. E. Brennan. 1991. Grounding in Communication. Perspectives on Socially Shared Cognition/American Psychological Association (1991), 127–149.
- Cohen et al. (1994) Sheldon Cohen, Tom Kamarck, Robin Mermelstein, et al. 1994. Perceived stress scale. Measuring stress: A guide for health and social scientists 10, 2 (1994), 1–2.
- Costa et al. (2016) Jean Costa, Alexander T Adams, Malte F Jung, François Guimbretière, and Tanzeem Choudhury. 2016. EmotionCheck: leveraging bodily signals and false feedback to regulate our emotions. In Proceedings of the 2016 ACM international joint conference on pervasive and ubiquitous computing. 758–769.
- Donnelly and Murray (1991) Daniel A Donnelly and Edward J Murray. 1991. Cognitive and emotional changes in written essays and therapy interviews. Journal of Social and Clinical psychology 10, 3 (1991), 334–350.
- Easwara Moorthy and Vu (2015) Aarthi Easwara Moorthy and Kim-Phuong L Vu. 2015. Privacy concerns for use of voice activated personal assistant in the public space. International Journal of Human-Computer Interaction 31, 4 (2015), 307–335.
- Enge (2020) Eric Enge. 2020. https://www.perficient.com/insights/research-hub/voice-usage-trends.
- Esterling et al. (1994) Brian A Esterling, Michael H Antoni, Mary Ann Fletcher, Scott Margulies, and Neil Schneiderman. 1994. Emotional disclosure through writing or speaking modulates latent Epstein-Barr virus antibody titers. Journal of consulting and clinical psychology 62, 1 (1994), 130.
- Esterling et al. (1999) Brian A Esterling, Luciano L’Abate, Edward J Murray, and James W Pennebaker. 1999. Empirical foundations for writing in prevention and psychotherapy: Mental and physical health outcomes. Clinical psychology review 19, 1 (1999), 79–96.
- Fadhil et al. (2018) Ahmed Fadhil, Gianluca Schiavo, Yunlong Wang, and Bereket A Yilma. 2018. The effect of emojis when interacting with conversational interface assisted health coaching system. In Proceedings of the 12th EAI international conference on pervasive computing technologies for healthcare. 378–383.
- Frisina et al. (2004) Pasquale G Frisina, Joan C Borod, and Stephen J Lepore. 2004. A meta-analysis of the effects of written emotional disclosure on the health outcomes of clinical populations. The Journal of nervous and mental disease 192, 9 (2004), 629–634.
- Grieger et al. (2021) Florian Grieger, Holger Klapperich, and Marc Hassenzahl. 2021. Trash it, punch it, burn it–using virtual reality to support coping with negative thoughts. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems. 1–6.
- Hammen (2005) Constance Hammen. 2005. Stress and depression. Annu. Rev. Clin. Psychol. 1 (2005), 293–319.
- Hardey (2002) Michael Hardey. 2002. ‘The story of my illness’: Personal accounts of illness on the Internet. Health: 6, 1 (2002), 31–46.
- Hart and Staveland (1988) Sandra G Hart and Lowell E Staveland. 1988. Development of NASA-TLX (Task Load Index): Results of empirical and theoretical research. In Advances in psychology. Vol. 52. Elsevier, 139–183.
- Howe et al. (2022) Esther Howe, Jina Suh, Mehrab Bin Morshed, Daniel McDuff, Kael Rowan, Javier Hernandez, Marah Ihab Abdin, Gonzalo Ramos, Tracy Tran, and Mary P Czerwinski. 2022. Design of Digital Workplace Stress-Reduction Intervention Systems: Effects of Intervention Type and Timing. In CHI Conference on Human Factors in Computing Systems. 1–16.
- Hutchby (2005) Ian Hutchby. 2005. ” Active Listening”: Formulations and the Elicitation of Feelings-Talk in Child Counselling. Research on language and Social Interaction 38, 3 (2005), 303–329.
- Iacovelli and Johnson (2012) Anthony M Iacovelli and Craig Johnson. 2012. Disclosure through face-to-face and instant messaging modalities: Psychological and physiological effects. Journal of Social and Clinical Psychology 31, 3 (2012), 225–250.
- Igarashi et al. (2022) Tasuku Igarashi, Shimpei Okuda, and Kazutoshi Sasahara. 2022. Development of the japanese version of the linguistic inquiry and word count dictionary 2015. Frontiers in psychology 13 (2022), 841534.
- Kim (2002) Heejung S Kim. 2002. We talk, therefore we think? A cultural analysis of the effect of talking on thinking. Journal of personality and social psychology 83, 4 (2002), 828.
- Kim et al. (2018) Songhyun Kim, Junseok Goh, and Soojin Jun. 2018. The use of voice input to induce human communication with banking chatbots. In Companion of the 2018 ACM/IEEE International Conference on Human-Robot Interaction. 151–152.
- Kocielnik et al. (2018) Rafal Kocielnik, Daniel Avrahami, Jennifer Marlow, Di Lu, and Gary Hsieh. 2018. Designing for workplace reflection: a chat and voice-based conversational agent. In Proceedings of the 2018 designing interactive systems conference. 881–894.
- Kudo (2005) Taku Kudo. 2005. Mecab: Yet another part-of-speech and morphological analyzer. http://mecab. sourceforge. net/ (2005).
- Lee et al. (2016) Sang Won Lee, Inyeop Kim, Jaehyun Yoo, Sungkyu Park, Bumseok Jeong, and Meeyoung Cha. 2016. Insights from an expressive writing intervention on Facebook to help alleviate depressive symptoms. Computers in Human Behavior 62 (2016), 613–619.
- Lee and Schwarz (2011) Spike WS Lee and Norbert Schwarz. 2011. Wiping the slate clean: Psychological consequences of physical cleansing. Current directions in psychological science 20, 5 (2011), 307–311.
- Lee and Schwarz (2021) Spike WS Lee and Norbert Schwarz. 2021. Grounded procedures: A proximate mechanism for the psychology of cleansing and other physical actions. Behavioral and Brain Sciences 44 (2021), e1.
- Li et al. (2022) Jinyu Li et al. 2022. Recent advances in end-to-end automatic speech recognition. APSIPA Transactions on Signal and Information Processing 11, 1 (2022).
- Luo (2022) Yuhan Luo. 2022. Promoting Rich and Low-Burden Self-Tracking With Multimodal Data Input. Ph. D. Dissertation.
- MacIsaac et al. (2022) Angela MacIsaac, Aislin R Mushquash, and Christine Wekerle. 2022. Writing Yourself Well: Dispositional Self-Reflection Moderates the Effect of a Smartphone App-Based Journaling Intervention on Psychological Wellbeing across Time. Behaviour Change (2022), 1–17.
- Majima et al. (2017) Yoshimasa Majima, Kaoru Nishiyama, Aki Nishihara, and Ryosuke Hata. 2017. Conducting online behavioral research using crowdsourcing services in Japan. Frontiers in Psychology 8 (2017), 378.
- McDuff et al. (2012) Daniel McDuff, Amy Karlson, Ashish Kapoor, Asta Roseway, and Mary Czerwinski. 2012. AffectAura: an intelligent system for emotional memory. In Proceedings of the SIGCHI conference on human factors in computing systems. 849–858.
- Murray et al. (1989) Edward J Murray, Alisa D Lamnin, and Charles S Carver. 1989. Emotional expression in written essays and psychotherapy. Journal of social and clinical psychology 8, 4 (1989), 414–429.
- Murray and Segal (1994) Edward J Murray and Daniel L Segal. 1994. Emotional processing in vocal and written expression of feelings about traumatic experiences. Journal of traumatic stress 7 (1994), 391–405.
- Mushtaq et al. (2018) Mamoona Mushtaq et al. 2018. Stress, headache, high blood pressure and self-esteem in working women. Pakistan Heart Journal 51, 2 (2018).
- Nussbaum (1995) Martha C Nussbaum. 1995. Objectification. Philosophy & Public Affairs 24, 4 (1995), 249–291.
- Park et al. (2021) SoHyun Park, Anja Thieme, Jeongyun Han, Sungwoo Lee, Wonjong Rhee, and Bongwon Suh. 2021. “I wrote as if I were telling a story to someone I knew.”: Designing Chatbot Interactions for Expressive Writing in Mental Health. In Designing Interactive Systems Conference 2021. 926–941.
- Peesapati et al. (2010) S Tejaswi Peesapati, Victoria Schwanda, Johnathon Schultz, Matt Lepage, So-yae Jeong, and Dan Cosley. 2010. Pensieve: supporting everyday reminiscence. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 2027–2036.
- Pennebaker (1997) James W Pennebaker. 1997. Writing about emotional experiences as a therapeutic process. Psychological science 8, 3 (1997), 162–166.
- Pennebaker and Beall (1986) James W Pennebaker and Sandra K Beall. 1986. Confronting a traumatic event: toward an understanding of inhibition and disease. Journal of abnormal psychology 95, 3 (1986), 274.
- Pennebaker and Chung (2007) James W Pennebaker and Cindy K Chung. 2007. Expressive writing, emotional upheavals, and health. Foundations of health psychology (2007), 263–284.
- Pennebaker et al. (2001) James W Pennebaker, Martha E Francis, and Roger J Booth. 2001. Linguistic inquiry and word count: LIWC 2001. Mahway: Lawrence Erlbaum Associates 71, 2001 (2001), 2001.
- Radford et al. (2022) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2022. Robust speech recognition via large-scale weak supervision. arXiv preprint arXiv:2212.04356 (2022).
- Reinhold et al. (2018) Maren Reinhold, Paul-Christian Bürkner, and Heinz Holling. 2018. Effects of expressive writing on depressive symptoms―A meta-analysis. Clinical Psychology: Science and Practice 25, 1 (2018), 80.
- Rogers and Farson (1957) Carl Ransom Rogers and Richard Evans Farson. 1957. Active listening. Industrial Relations Center of the University of Chicago Chicago, IL.
- Ruan et al. (2018) Sherry Ruan, Jacob O Wobbrock, Kenny Liou, Andrew Ng, and James A Landay. 2018. Comparing speech and keyboard text entry for short messages in two languages on touchscreen phones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 4 (2018), 1–23.
- Rzepka et al. (2022) Christine Rzepka, Benedikt Berger, and Thomas Hess. 2022. Voice assistant vs. Chatbot–examining the fit between conversational agents’ interaction modalities and information search tasks. Information Systems Frontiers 24, 3 (2022), 839–856.
- Sas et al. (2016) Corina Sas, Steve Whittaker, and John Zimmerman. 2016. Design for rituals of letting go: An embodiment perspective on disposal practices informed by grief therapy. ACM Transactions on Computer-Human Interaction (TOCHI) 23, 4 (2016), 1–37.
- Sauter (2014) Theresa Sauter. 2014. ‘What’s on your mind?’Writing on Facebook as a tool for self-formation. New media & society 16, 5 (2014), 823–839.
- Sloan and Marx (2004) Denise M Sloan and Brian P Marx. 2004. Taking pen to hand: evaluating theories underlying the written disclosure paradigm. Clinical psychology: Science and practice 11, 2 (2004), 121.
- Soesilo et al. (2021) Primidya KM Soesilo, Maureen L Morrin, and Nese Nur Yazgan Onuklu. 2021. No longer green with envy: Objectifying and destroying negative consumer emotions. Journal of Consumer Affairs 55, 3 (2021), 1111–1138.
- Takami et al. (2022) Taiki Takami, Shohei Ando, Shota Nakayama, Asahi Saito, Yui Suga, Takumi Hamazaki, Yutaro Yano, and Hiroyuki Kajimoto. 2022. StressMincer: Enhancement of Catharsis Effect by Visualization of Words and Destruction with Haptic Feedback. In Proceedings of the Augmented Humans International Conference 2022. 344–346.
- Tong et al. (2023) Xin Tong, Matthew Louis Mauriello, Marco Antonio Mora-Mendoza, Nina Prabhu, Jane Paik Kim, and Pablo E Paredes Castro. 2023. Just Do Something: Comparing Self-proposed and Machine-recommended Stress Interventions among Online Workers with Home Sweet Office. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–20.
- Treem and Leonardi (2013) Jeffrey W Treem and Paul M Leonardi. 2013. Social media use in organizations: Exploring the affordances of visibility, editability, persistence, and association. Annals of the International Communication Association 36, 1 (2013), 143–189.
- Veit and Ware (1983) Clairice T Veit and John E Ware. 1983. The structure of psychological distress and well-being in general populations. Journal of consulting and clinical psychology 51, 5 (1983), 730.
- Walther (1993) Joseph B Walther. 1993. Impression development in computer-mediated interaction. Western Journal of Communication (includes Communication Reports) 57, 4 (1993), 381–398.
- Wang et al. (2018) Liuping Wang, Xiangmin Fan, Feng Tian, Lingjia Deng, Shuai Ma, Jin Huang, and Hongan Wang. 2018. Mirroru: Scaffolding emotional reflection via in-situ assessment and interactive feedback. In Extended Abstracts of the 2018 CHI Conference on Human Factors in Computing Systems. 1–6.
- Wright and Chung (2001) Jeannie Wright and Man Cheung Chung. 2001. Mastery or mystery? Therapeutic writing: A review of the literature. British Journal of Guidance & Counselling 29, 3 (2001), 277–291.
- Wu and Bailey (2018) Y Wayne Wu and Brian P Bailey. 2018. Soften the pain, increase the gain: Enhancing users’ resilience to negative valence feedback. Proceedings of the ACM on Human-Computer Interaction 2, CSCW (2018), 1–20.
- Xu et al. (2012) Alison Jing Xu, Rami Zwick, and Norbert Schwarz. 2012. Washing away your (good or bad) luck: Physical cleansing affects risk-taking behavior. Journal of Experimental Psychology: General 141, 1 (2012), 26.
- Yu et al. (2019) Qian Yu, Tonya Nguyen, Soravis Prakkamakul, and Niloufar Salehi. 2019. ” I almost fell in love with a machine” speaking with computers affects self-disclosure. In Extended Abstracts of the 2019 CHI conference on human factors in computing systems. 1–6.
- Zhong and Liljenquist (2006) Chen-Bo Zhong and Katie Liljenquist. 2006. Washing away your sins: Threatened morality and physical cleansing. Science 313, 5792 (2006), 1451–1452.
Appendix A Questionnaires Used in Study 1
The original questionnaire was in Japanese, and we translated it into English for the report in this paper.
| Pre-experiment Questionnaire | |
| Demographic Information | |
| Please provide your username on the crowdsourcing website. | |
| Please specify your gender. | |
| Please specify your age group. | |
| PSS questions | |
| 14 questions from PSS about the past week. | |
| Chatbot registration and confirmation | |
| Did you register the chatbot from the link above and send your username on the crowdsourcing website? (Yes) | |
| Post-experiment Questionnaire | |
| PSS questions | |
| 14 questions from PSS about the past week. | |
| Overall experience of the system | |
| How easy was it to use the system? (1–5) | |
| Do you think providing descriptions of your stress helped in relieving stress? (1–5) | |
| Please explain why you answered the above question as you did. | |
| Keyboard-based text entry | |
| Was describing your stress using keyboard-based text input easy to do? (1–5) | |
| Do you believe that keyboard-based text input helped in relieving stress? (1–5) | |
| Please share your thoughts and feedback on keyboard-based text input. | |
| Workload (NASA-TLX) with keyboard-based text input. | |
| Voice messages | |
| Was describing your stress using voice messages easy to do? (1–5) | |
| Do you believe that voice messages helped in relieving stress? (1–5) | |
| Please share your thoughts and feedback on voice messages. | |
| Workload (NASA-TLX) with voice messages. | |
| Other questions | |
| Which input method do you prefer? | |
| Why do you prefer that input method? | |
| Do you use voice input regularly in your daily life? (1–5) | |
| Were you aware that the primary purpose of this survey was an experiment on input methods? (1–5) | |
| If you have any opinions or advice regarding stress relief through the system, please share. | |
| Is there anything else you would like to add? | |
| In the future, we may request your participation in additional surveys related to this research. May we contact you for such purposes? (Yes/No) | |
Appendix B Questionnaires Used in Study 2
The original questionnaire was in Japanese, and we translated it into English for the report in this paper.
| Pre-experiment Questionnaire | |
| Demographic Information | |
| Please provide your username on the crowdsourcing website. | |
| Please specify your gender. | |
| Please specify your age group. | |
| PSS questions | |
| 14 questions from PSS about the past week. | |
| Chatbot registration and confirmation | |
| Did you register the chatbot from the link above and send your username on the crowdsourcing website? (Yes) | |
| Post-experiment Questionnaire | |
| PSS questions | |
| 14 questions from PSS about the past week. | |
| Experience of None Feedback | |
| 12 questions (in Table 5) about None feedback. | |
| Workload during None Feedback: NASA-TLX. | |
| Experience of Passive Feedback | |
| 12 questions (in Table 5) about Passive feedback. | |
| Please share any preferences for other types of videos you would like to see. | |
| Workload during Passive Feedback: NASA-TLX. | |
| Experience of Interactive Feedback | |
| 12 questions (in Table 5) about Interactive feedback. | |
| Please share any preferences for other types of interactions. | |
| Workload during Interactive Feedback: NASA-TLX. | |
| Other questions | |
| Where did you primarily use the chatbot? (Home/Work/Commute/Other) | |
| How do you usually cope with stress in your daily life? (Rest, Relaxation/Exercise/Talking or meeting with family and friends/Eating favorite foods/Playing games/Drawing or playing music/Shopping or going to karaoke/Other) | |
| Which type of feedback do you prefer to use? (None/Passive/Interactive) | |
| Why do you prefer that type of feedback? | |
| Do you wish to continue using this chatbot? (1: Not at all–5: Strongly Agree) | |
| On a scale of 0–10, how likely are you to recommend this chatbot to friends or colleagues? | |
| What other features would you like to see in addition to video replies and website usage? | |
| Please share any opinions or advice regarding stress relief through using the chatbot. | |
| Is there anything else you would like to add? | |
| In the future, we may request your participation in additional surveys related to this research. May we contact you for such purposes? (Yes/No) | |
B.1. 12 Questions about Each Feedback Design
| Eight 5-point Likert scale questions (1: Not at all–5: Strongly Agree) | |
| This feedback was easy to use. | |
| This feedback helps in reducing stress. | |
| This feedback makes you more inclined to provide descriptions. | |
| This feedback makes you more likely to use our system. | |
| This feedback requires effort. | |
| You are satisfied with This Feedback. | |
| You intend to continue using This Feedback. | |
| You have complaints about This Feedback. | |
| Four open-ended questions | |
| When would you like to use This Feedback (or not use it)? | |
| Please share the positive aspects of This Feedback. | |
| Please share the negative aspects of This Feedback. | |
| Any other opinions or feedback regarding This Feedback. | |